# Motivic cohomology of cyclic coverings

### v2

@@ 3 Cyclic coverings and their motivic cohomology
Assumption (b) and Remark 3.16 guarantee that the Galois action of $\mu_{s}(k)$ is trivial on $H^{p,q}(Y_{s}\setminus F,R)$ as soon as the projection $\pi:(Y_{s}\setminus F)\times_{k}\mathbb{G}_{m,k}\rightarrow Y_{s}\setminus F$ induces an isomorphism
@@
The reasoning in the proofs of Theorems 3.19 and 3.20 also work for Chow groups tensored with a coefficient ring $R$ such that $s\in R^{\times}$:
@@
We can now prove several corollaries to Theorems 3.19, 3.20 and Theorem 3.21:
@@
One can assume $p\leq 2q$. Then the statement follows inductively from Theorem 3.20, Theorem 3.19 and the fact that $H^{p,q}(Spec(k),R)=0$ for $p\neq 0$ (cf. [MVW, Preface]): Note that, for each $(p,q)\in\mathbb{Z}^{2}$ with $p\leq 2q$, the base case is bidegree $(2p-2q,p-q)\in\mathbb{Z}^{2}$ (which is covered by Theorem 3.20) and applying Theorem 3.19 exactly $2q-p$ times yields the claim for bidegree $(p,q)$. ∎
@@
Under the assumptions of Corollary 3.26, further assume that $n=2$ and $\dim(Y_{s})=4$. Then $H_{Nis}^{2}(Y_{s},\textbf{I}^{3})=0$.
@@
$\rightarrow H^{2+j,j-1}(Y_{s},\mathbb{Z}/2\mathbb{Z})\rightarrow H^{2+j,j}(Y_{s},\mathbb{Z}/2\mathbb{Z})\rightarrow H^{2}_{Nis}(Y_{s},\overline{\textbf{I}}^{j})\rightarrow H^{3+j,j-1}(Y_{s},\mathbb{Z}/2\mathbb{Z})\rightarrow...$
Now it follows directly from Corollary 3.26 that $H^{2}_{Nis}(Y_{s},\overline{\textbf{I}}^{3})=H^{2}_{Nis}(Y_{s},\overline{\textbf{I}}^{4})=0$. Note that $\textbf{I}^{4}=\overline{\textbf{I}}^{4}$ over $Y_{s}$ by [AF1, Proposition 5.1]. The statement now follows from the long exact sequence of cohomology groups associated to the short exact sequence of sheaves
@@
over the small Nisnevich site of $Y_{s}$. ∎
Remark 3.29.
By inspection of the proof of Corollary 3.28, one sees easily that it is sufficient the replace the assumptions (c) and (d) from Theorem 3.26 by assumptions (c’) and (d’) of Remark 3.27.
@@
[…]
@@
[…]
@@
Let $s,t\geq 1$ be integers. Under the assumptions introduced at the beginning of this section, we call the subvariety $Y_{s,t}$ of $X\times_{k}\mathbb{A}^{1}_{k}\times_{k}\mathbb{A}^{1}_{k}$ defined by the equation $\{f+u^{s}+v^{t}=0\}$ a bicyclic covering of orders $s,t$ (here $u$ and $v$ are the variables of the two copies of $\mathbb{A}^{1}_{k}$). We conclude the section with some results on the Chow groups and on the vector bundles of such bicyclic coverings. At first, we prove a result on the Chow groups of such bicyclic coverings:
@@
With the notation and the assumptions from the beginning of the section, further assume that
@@
The variety $Y_{s,t}$ is a cyclic covering of the product of $X$ and $\mathbb{A}^{1}_{\mathbb{C}}$ (with coordinate $u$) along $Y_{s}=\{-f-u^{s}=0\}$ of order $t$; note that $Y_{s}$ is a cyclic covering of $X$ branched to order $s$ as usual, but here we use the regular function $-f$ instead of $f$. Theorem 3.21 together with assumptions (d), (e) and Remark 3.17 implies that $CH^{i}(Y_{s,t})$, $i\geq 1$, are divisible by any integer $n$ with $\langle n,t\rangle=\mathbb{Z}$ (just take $R=\mathbb{Z}/n\mathbb{Z}$). Now realize that $Y_{s,t}$ is also a cyclic covering of the product of $X$ and $\mathbb{A}^{1}_{\mathbb{C}}$ (with coordinate $v$) along $Y_{t}=\{-f-v^{t}=0\}$ of order $s$. Using Theorem 3.21 again yields that $CH^{i}(Y_{s,t})$, $i\geq 1$, are also divisible by any integer $n$ with $\langle n,s\rangle=\mathbb{Z}$ (again, just take $R=\mathbb{Z}/n\mathbb{Z}$). As $s$ and $t$ are coprime, the groups $CH^{i}(Y_{s,t})$, $i\geq 1$, are divisible by any prime number and hence are divisible groups. In conclusion, it suffices to show that there is an integer $n\neq 0$ such that $CH^{i}(Y_{s,t})$ is $n$-torsion.
@@
Finally, we can give a sufficient criterion for the triviality of all vector bundles over a bicyclic covering as in Theorem 3.32 if the dimension of the bicyclic covering is $4$:
@@
With the notation and the assumptions from the beginning of the section, further assume that
@@
    the smooth variety $X$ is $\mathbb{A}^{1}$-contractible
@@
    the smooth variety $F_{0}=\{f=0\}\subset X$ is $\mathbb{A}^{1}$-contractible
@@
    $\langle d,s\rangle=\langle d,t\rangle=\langle s,t\rangle=\mathbb{Z}$ and both $s$ and $t$ are odd
@@
Since $\dim(Y_{s,t})=4$, the triviality of all stably trivial vector bundles of rank $2$ over $Y_{s,t}$ is equivalent to the triviality of $W_{SL}(Y_{s,t})$ (cf. [Sy, Theorem 3.19]). So it suffices to prove the first statement in the theorem. The variety $Y_{s,t}$ has trivial Chow groups by Theorem 3.32. In particular, $Y_{s,t}$ has only stably trivial vector bundles. Since $\dim(Y_{s,t})=4$, it now suffices by Theorem 2.26 to show that $H^{2}_{Nis}(Y_{s,t},\textbf{I}^{3})$ is trivial. We can first apply Corollary 3.26 in order to deduce that $H^{p,q}(Y_{s},\mathbb{Z}/2\mathbb{Z})=0$ for any $(p,q)\in\mathbb{Z}^{2}$ with $p>q\geq 1$. Then we can apply Corollary 3.28 (and Remark 3.29) to obtain that indeed $H^{2}_{Nis}(Y_{s,t},\textbf{I}^{3})=0$. This finishes the proof. ∎
@@ 4 Examples
Finally, we are able to discuss some examples of smooth affine varieties which are cyclic coverings of varieties considered in Example 2.16 and of varieties considered in [DPØ]. As usual, we assume that the base field $k$ is algebraically closed of characteristic 00.
@@ 4.1 Cyclic coverings of Koras-Russell threefolds of the first kind
with $f=y\in\mathcal{O}_{X}(X)$ and $s=\alpha_{1}$ (see Example 2.16). There is a $\mathbb{G}_{m,k}$-action on $X$ given by $(\lambda,x,y,z,t)\mapsto(\lambda^{\alpha_{2}\alpha_{3}}x,\lambda^{-\alpha_{2}\alpha_{3}}y,\lambda^{\alpha_{3}}z,\lambda^{\alpha_{2}}t)$ which makes $f=y$ a quasi-invariant of weight $-\alpha_{2}\alpha_{3}$.
@@
If $k=\mathbb{C}$, then $Y_{s}$ is topologically contractible (it is a Koras-Russell threefold by [Z, Proposition 6.2]); it is well-known that topologically contractible smooth affine complex varieties have a trivial Picard group and hence $CH^{1}(Y_{s})\cong CH^{1}(Y_{s})\otimes_{\mathbb{Z}}R=0$ (cf. [G, Theorem 1]). If $\alpha_{1},\alpha_{2},\alpha_{3}>>1$, then [Z, Proposition 6.2] shows that the logarithmic Kodaira dimension of $Y_{s}$ is $2$ and therefore the variety $Y_{s}$ has to be a Koras-Russell threefold of the third kind in this case as Koras-Russell threefolds of the first or second kind are known to have logarithmic Kodaira dimension $\infty$- ∞ (see Example 2.16).
@@ 4.2 Cyclic coverings of Koras-Russell threefolds of the second kind
If $k=\mathbb{C}$, it follows directly from the construction in [KR, Theorem 4.1] that the variety $Y_{s}$ is a Koras-Russell threefold; indeed, choose $-a^{\prime}_{1}=dl-1$, $a^{\prime}_{2}=d$, $a^{\prime}_{3}=1$, $\bar{u}_{2}=\xi$, $\bar{u}_{3}=\zeta+{(\zeta^{d}+\xi)}^{l}$ in [KR, Theorem 4.1] in order to obtain the type of equation defining $Y_{s}$. In particular, $Y_{s}$ is topologically contractible and has a trivial Picard group, i.e., $CH^{1}(Y_{s})\cong CH^{1}(Y_{s})\otimes_{\mathbb{Z}}R=0$. If $\alpha_{2},\alpha_{3}>>1$ and $\alpha_{1}\geq(dl-1)\alpha_{2}\alpha_{3}$, this Koras-Russell threefold has logarithmic Kodaira dimension $2$ by [KR, Proposition 6.5] and hence must be a Koras-Russell threefold of the third kind.

### v3

@@ 3 Cyclic coverings and their motivic cohomology
Assumption (b) and Lemma 3.14 guarantee that the Galois action of $\mu_{s}(k)$ is trivial on $H^{p,q}(Y_{s}\setminus F,R)$ as soon as the projection $\pi:(Y_{s}\setminus F)\times_{k}\mathbb{G}_{m,k}\rightarrow Y_{s}\setminus F$ induces an isomorphism
@@
The reasoning in the proofs of Theorems 3.17 and 3.18 also work for Chow groups tensored with a coefficient ring $R$ such that $s\in R^{\times}$:
@@
Example 3.20.
A simple example satisfying the hypotheses of Theorem 3.19 is $X=\mathbb{A}^{n}_{k}=Spec(k[x_{1},...,x_{n}])$ with $\mathbb{G}_{m,k}$-action given by $(\lambda,x_{1},...,x_{n})\mapsto(x_{1}\lambda^{d_{1}},...,x_{n}\lambda^{d_{n}})$ for some $d_{j}\in\mathbb{Z}$. In this case any coordinate function $x_{j}$ is a quasi-invariant of weight $d_{j}$ and, if we fix some $j\in\{1,...,n\}$, the cyclic covering $Y_{s}$ of $X$ of order $s$ with respect to the coordinate function $x_{j}$ is again isomorphic to an affine space of dimension $n$ over $k$. If $s\in R^{\times}$ and $\langle d_{j},s\rangle=\mathbb{Z}$, then Theorem 3.19 implies that $CH^{i}(X)\otimes_{\mathbb{Z}}R\cong CH^{i}(Y_{s})\otimes_{\mathbb{Z}}R$ for $i\geq 0$ (which is clear as both $X$ and $Y_{s}$ are affine spaces of dimension $n$ over $k$).
Non-trivial examples satisfying the hypotheses of Theorem 3.19 are given in Section 4.
@@
We can now prove several corollaries to Theorems 3.17, 3.18 and Theorem 3.19:
@@
One can assume $p\leq 2q$. Then the statement follows inductively from Theorem 3.18, Theorem 3.17 and the fact that $H^{p,q}(Spec(k),R)=0$ for $p\neq 0$ (cf. [MVW, Preface]): Note that, for each $(p,q)\in\mathbb{Z}^{2}$ with $p\leq 2q$, the base case is bidegree $(2p-2q,p-q)\in\mathbb{Z}^{2}$ (which is covered by Theorem 3.18) and applying Theorem 3.17 exactly 2qp2q-p times yields the claim for bidegree $(p,q)$. ∎
@@
Under the assumptions of Corollary 3.24 for $R=\mathbb{Z}/2\mathbb{Z}$, further assume that $\dim(Y_{s})=4$. Then $H^{2}_{Nis}(X,\textbf{I}^{3})\cong H_{Nis}^{2}(Y_{s},\textbf{I}^{3})$.
@@
$\rightarrow H^{2+j,j-1}(V,\mathbb{Z}/2\mathbb{Z})\rightarrow H^{2+j,j}(V,\mathbb{Z}/2\mathbb{Z})\rightarrow H^{2}_{Nis}(V,\overline{\textbf{I}}^{j})\rightarrow H^{3+j,j-1}(V,\mathbb{Z}/2\mathbb{Z})\rightarrow...$
for any smooth affine fourfold $V$ over $k$. Now [MVW, Vanishing Theorem 19.3] directly implies that the long exact sequence above for $j=3$ yields an isomorphism $H^{5,3}(V,\mathbb{Z}/2\mathbb{Z})\cong H^{2}_{Nis}(V,\overline{\textbf{I}}^{3})$ for any smooth affine fourfold $V$ over $k$.
Note that $\textbf{I}^{4}=\overline{\textbf{I}}^{4}$ over $V$ by [AF1, Proposition 5.1], so the proof of [F, Lemma 4.0.3] shows that $H^{2}_{Nis}(V,\textbf{I}^{4})\cong H^{2}_{Nis}(V,\overline{\textbf{I}}^{4})\cong H^{2}_{Nis}(V,\textbf{K}_{4}^{M}/2)=0$ as well as $H^{3}_{Nis}(V,\textbf{I}^{4})\cong H^{3}_{Nis}(V,\overline{\textbf{I}}^{4})\cong H^{3}_{Nis}(V,\textbf{K}_{4}^{M}/2)=0$. Now the long exact sequence of cohomology groups associated to the short exact sequence of sheaves
@@
over the small Nisnevich site of $V$ shows that $H^{2}_{Nis}(V,\textbf{I}^{3})\cong H^{2}_{Nis}(V,\overline{\textbf{I}}^{3})$ and, in particular, $H^{5,3}(V,\mathbb{Z}/2\mathbb{Z})\cong H^{2}_{Nis}(V,\textbf{I}^{3})$ by the first paragraph of this proof.
Finally, as both $X$ and $Y_{s}$ are smooth affine fourfolds over $k$, Corollary 3.24 shows that $\varphi_{s}$ induces an isomorphism $H^{5,3}(X,\mathbb{Z}/2\mathbb{Z})\cong H^{5,3}(Y_{s},\mathbb{Z}/2\mathbb{Z})$ and therefore we obtain an isomorphism $H^{2}_{Nis}(X,\textbf{I}^{3})\cong H_{Nis}^{2}(Y_{s},\textbf{I}^{3})$. This finishes the proof. ∎
@@
${H^{p,q}(X,R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{id}_{X}^{\ast}}$$\scriptstyle{\varphi_{s}^{\ast}}\scriptstyle{\cong}$${H^{p,q}(Y_{s},R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g^{\ast}}\scriptstyle{\cong}$${H^{p,q}(X,R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{s}^{\ast}}\scriptstyle{\cong}$${H^{p,q}(Y_{s},R)}$
@@
${H^{p,q}(Y_{s},R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i^{\ast}}\scriptstyle{\cong}$$\scriptstyle{g^{\ast}}$${H^{p,q}(Y_{s},R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i^{\ast}}\scriptstyle{\cong}$${H^{p,q}(Y_{s}\setminus F,R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g^{\ast}}$${H^{p,q}(Y_{s}\setminus F,R)}$
@@
Let $s,t\geq 1$ be integers. With the notation of Definition 1.1 and under the General Assumptions 1.2, we call the subvariety $Y_{s,t}$ of $X\times_{k}\mathbb{A}^{1}_{k}\times_{k}\mathbb{A}^{1}_{k}$ defined by the equation $\{f+u^{s}+v^{t}=0\}$ a bicyclic covering of orders $s,t$ (here $u$ and $v$ are the variables of the two copies of $\mathbb{A}^{1}_{k}$). We conclude the section with some results on the Chow groups and on the vector bundles of such bicyclic coverings. At first, we prove a result on the Chow groups of such bicyclic coverings:
@@
With the notation of Definition 1.1 and under the General Assumptions 1.2, further assume that
@@
The variety $Y_{s,t}$ is a cyclic covering of the product of $X$ and $\mathbb{A}^{1}_{k}$ (with coordinate $u$) along $Y_{s}=\{-f-u^{s}=0\}$ of order $t$; note that $Y_{s}$ is a cyclic covering of $X$ branched to order $s$ as usual, but here we use the regular function f-f instead of $f$. Theorem 3.19 together with assumptions (d), (e) and Remark 3.15 implies that $CH^{i}(Y_{s,t})$, $i\geq 1$, are divisible by any integer $n$ with $\langle n,t\rangle=\mathbb{Z}$ (just take $R=\mathbb{Z}/n\mathbb{Z}$). Now realize that $Y_{s,t}$ is also a cyclic covering of the product of $X$ and $\mathbb{A}^{1}_{k}$ (with coordinate $v$) along $Y_{t}=\{-f-v^{t}=0\}$ of order $s$. Using Theorem 3.19 again yields that $CH^{i}(Y_{s,t})$, $i\geq 1$, are also divisible by any integer $n$ with $\langle n,s\rangle=\mathbb{Z}$ (again, just take $R=\mathbb{Z}/n\mathbb{Z}$). As $s$ and $t$ are coprime, the groups $CH^{i}(Y_{s,t})$, $i\geq 1$, are divisible by any prime number and hence are divisible groups. In conclusion, it suffices to show that there is an integer $n\neq 0$ such that $CH^{i}(Y_{s,t})$ is $n$-torsion.
@@
Example 3.32.
Again, a simple example is $X=\mathbb{A}^{n}_{k}=Spec(k[x_{1},...,x_{n}])$ with $\mathbb{G}_{m,k}$-action given by $(\lambda,x_{1},...,x_{n})\mapsto(x_{1}\lambda^{d_{1}},...,x_{n}\lambda^{d_{n}})$ for some $d_{j}\in\mathbb{Z}$. As explained before, any coordinate function $x_{j}$ is then a quasi-invariant of weight $d_{j}$. If we then fix some $j\in\{1,...,n\}$, the bicyclic covering $Y_{s,t}$ of $X$ of orders $s,t$ with respect to the coordinate function $x_{j}$ is isomorphic to an affine space of dimension $n+1$ over $k$. If $s,t$ and $d_{j}$ are pairwise coprime, then Theorem 3.30 recovers the fact that $CH^{i}(Y_{s,t})=0$ for $i\geq 1$.
A non-trivial example satisfying the hypotheses of Theorem 3.30 is given in Section 4.4. Finally, we can give a sufficient criterion for the triviality of all vector bundles over a bicyclic covering as in Theorem 3.30 if the dimension of the bicyclic covering is $4$:
@@
With the notation of Definition 1.1 and the General Assumptions 1.2, further assume that
@@
    the smooth variety $X$ has the Chow groups of $Spec(k)$
@@
    the smooth variety $F_{0}=\{f=0\}\subset X$ has the Chow groups of $Spec(k)$
@@
Since $\dim(Y_{s,t})=4$, the triviality of all stably trivial vector bundles of rank $2$ over $Y_{s,t}$ is equivalent to the triviality of $W_{SL}(Y_{s,t})$ (cf. [Sy, Theorem 3.19]). So it suffices to prove the first statement in the theorem. The variety $Y_{s,t}$ has trivial Chow groups by Theorem 3.30. In particular, $Y_{s,t}$ has only stably trivial vector bundles. Since $\dim(Y_{s,t})=4$, it now suffices by Theorem 2.26 to show that $H^{2}_{Nis}(Y_{s,t},\textbf{I}^{3})$ is trivial.
Note that either $s$ or $t$ has to be odd, so $Y_{s,t}$ is a cyclic covering of $X\times_{k}\mathbb{A}^{1}_{k}$ of odd order in any case. As $H^{2i,i}(X,\mathbb{Z}/2\mathbb{Z})\cong CH^{i}(X)\otimes_{\mathbb{Z}}\mathbb{Z}/2\mathbb{Z}=0$ for $i\geq 1$ by assumption, Corollary 3.27 therefore implies that $H_{Nis}^{2}(Y_{s,t},\textbf{I}^{3})\cong H_{Nis}^{2}(X\times_{k}\mathbb{A}^{1}_{k},\textbf{I}^{3})$. But $H_{Nis}^{2}(X\times_{k}\mathbb{A}^{1}_{k},\textbf{I}^{3})\cong H_{Nis}^{2}(X,\textbf{I}^{3})=0$, where the first isomorphism follows from the fact that $\textbf{I}^{3}$ is a strictly $\mathbb{A}^{1}$-invariant sheaf of abelian groups (cf. [AF1, Section 2]) and the triviality of $H_{Nis}^{2}(X,\textbf{I}^{3})$ follows from [AF1, Proposition 5.2] as $X$ is a smooth affine threefold over $k$. This finishes the proof. ∎
Example 3.35.
If we let $n=3$ in Example 3.32, then Theorem 3.33 recovers the fact that all vector bundles over $\mathbb{A}^{4}_{k}$ are trivial.
A non-trivial example satisfying the hypotheses of Theorem 3.33 is given in Section 4.4.
@@ 4 Examples
Finally, we are able to discuss some examples of smooth affine varieties which are cyclic coverings of varieties considered in Example 2.16 and of varieties considered in [DPØ]. As usual, we assume that the base field $k$ is algebraically closed of characteristic 0.
@@ 4.1 Cyclic coverings of Koras-Russell threefolds of the first kind
with $f=y\in\mathcal{O}_{X}(X)$ and $s=\alpha_{1}$ (see Example 2.16). There is a $\mathbb{G}_{m,k}$-action on $X$ given by $(\lambda,x,y,z,t)\mapsto(\lambda^{\alpha_{2}\alpha_{3}}x,\lambda^{-\alpha_{2}\alpha_{3}}y,\lambda^{\alpha_{3}}z,\lambda^{\alpha_{2}}t)$ which makes $f=y$ a quasi-invariant of weight $\alpha_{2}\alpha_{3}$.
@@
If $k=\mathbb{C}$, then $Y_{s}$ is topologically contractible (it is a Koras-Russell threefold by [Z, Proposition 6.2]); it is well-known that topologically contractible smooth affine complex varieties have a trivial Picard group and hence $CH^{1}(Y_{s})\cong CH^{1}(Y_{s})\otimes_{\mathbb{Z}}R=0$ (cf. [G, Theorem 1]). If $\alpha_{1},\alpha_{2},\alpha_{3}>>1$, then [Z, Proposition 6.2] shows that the logarithmic Kodaira dimension of $Y_{s}$ is $2$ and therefore the variety $Y_{s}$ has to be a Koras-Russell threefold of the third kind in this case as Koras-Russell threefolds of the first or second kind are known to have logarithmic Kodaira dimension -$\infty$ (see Example 2.16).
@@ 4.2 Cyclic coverings of Koras-Russell threefolds of the second kind
If $k=\mathbb{C}$, it follows directly from the construction in [KR, Theorem 4.1] that the variety $Y_{s}$ is a Koras-Russell threefold; indeed, choose $a^{\prime}_{1}=dl-1$, $a^{\prime}_{2}=d$, $a^{\prime}_{3}=1$, $\bar{u}_{2}=\xi$, $\bar{u}_{3}=\zeta+{(\zeta^{d}+\xi)}^{l}$ in [KR, Theorem 4.1] in order to obtain the type of equation defining $Y_{s}$. In particular, $Y_{s}$ is topologically contractible and has a trivial Picard group, i.e., $CH^{1}(Y_{s})\cong CH^{1}(Y_{s})\otimes_{\mathbb{Z}}R=0$. If $\alpha_{2},\alpha_{3}>>1$ and $\alpha_{1}\geq(dl-1)\alpha_{2}\alpha_{3}$, this Koras-Russell threefold has logarithmic Kodaira dimension $2$ by [KR, Proposition 6.5] and hence must be a Koras-Russell threefold of the third kind.